# Domain Generalization: A Tale of Two ERMs

### v1

@@ C.4 Proof of Theorem˜2
$\mathbb{E}_{X,D,M}\left[\gamma(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\leq R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
@@
$f^{*}_{\mathrm{full}}(X,D)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XMD}\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DG}}.$
@@
$R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}$ $=\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{full}}(X,D)}\right]-\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$=\mathbb{E}_{X,D,M}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(X,D)\right)|X,D,M\right]$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d,M=m\right)$
$=\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d\right)$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d,M=m\right)$
@@
$f^{*}_{\mathrm{full}}(x,d)=f^{*}_{\mathrm{DG}}(x,m)\quad\text{ almost surely w.r.t. }P_{XMD}.\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DG}}$
C.5 Proof of Proposition˜3
Proposition  For function classes $\{\mathcal{F}_{k}\}$ and $\{\mathcal{G}_{k}\}$ satisfying Assumption 1. Under covariate shift,
$R_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\geq R_{\mathrm{DG},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f),\quad\forall k,$ (6)
@@ C.5 Proof of Proposition˜3
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R_{\mathrm{DG},\mathcal{F}_{k}}$ (7)
Proof. The first line follows from the first assumption of of Assumption 1, where every function $g\in\mathcal{G}_{k}$ is also realizable in $\mathcal{F}_{k}$.
The second line follows from covariate shift assumption
$R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$
the universal approximation property of function class $\mathcal{G}_{k}$
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=R^{*}_{\mathrm{pool}},$
and the fact that
$R^{*}_{\mathrm{DG}}\leq R_{\mathrm{DG},\mathcal{F}_{k}}\leq R_{\mathrm{pool},\mathcal{G}_{k}}.$
@@ Appendix D Experimental details
The following subsections follows the same structure as Section˜6, while providing additional details and full tables.
@@ D.1 Annotator disagreement
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate the sentence $x$ with the annotator profile $m$ using the text prompt shown in Figure˜3.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 10 epochs with learning rate $5\cdot 10^{-5}$. Table˜5 reports the results over three trials. Our models consistently outperform prior work, with the best configuration achieving over $90\%$ test accuracy—substantially higher than the previous state-of-the-art reported by Deng et al. (2023).
@@
Section˜D.1 presents the results. DI-ERM consistently outperforms pooling ERM across different feature extractors.
Table 6: Test accuracy on the sentiment disagreement dataset (frozen feature extractor). DI-ERM consistently outperforms pooling ERM, and in some settings surpasses the prior state-of-the-art of Deng et al. (2023). We highlight the best performance reported by Deng et al. (2023) (69.77) and our highest score (83.41). $\dagger$: Checkpoints used in Deng et al. (2023) were not publicly specified.
@@
AnnEmb (Deng et al., 2023) BERT 64.61
@@
Pooling ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 45.85
@@
gray!80blackPooling ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 55.07
@@
Pooling ERM (linear) bert-base-multilingual-uncased-sentiment 43.06
@@
gray!80blackPooling ERM (MLP) bert-base-multilingual-uncased-sentiment 53.90
@@ D.2 Reviewer-specific analysis
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate each review $x$ with reviewer context $m$, represented by 20 randomly selected reviews from the same reviewer, using the text prompt in Figure˜4.
We choose nomic-embed-text-v1.5 (Nussbaum et al., 2025), which supports a context window of up to 8012 tokens (we choose 2048 in our experiments), in order to handle the long reviwer context $m$.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 2 epochs with learning rate $10^{-5}$. Table˜7 reports the results over three trials. DI-ERM achieves the best performance, outperforming previously reported methods on the WILDS leaderboard (https://wilds.stanford.edu/).
@@
Table 7: Reviewer-specific sentiment analysis. DI-ERM (ours) achieves the highest accuracy, outperforming prior state-of-the-art by Yao et al. (2022).
@@
ERM nomic-embed-text 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
@@
Inspired by Zaheer et al. (2017); Deshmukh et al. (2018), we implement a Domain2Vec-style module to encode reviewer-specific domain information. Given a set of reviews $\{x_{1},x_{2},\dots,x_{n}\}\sim P_{X|D=d}$ written by reviewer $d$, we learn a mapping
@@
Section˜D.2 shows the result.
Table 8: Sentiment classification on Amazon-WILDS with reviewer-specific signals. “Domain2Vec” denotes reviewer encoding based on a learned mean embedding. DI-ERM variants consistently outperform pooling ERM baselines.
@@
Pooling ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 67.42 48.00
@@
gray!80blackPooling ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 67.59 48.00
@@
gray!80blackDI-ERM (Domain2Vec) distilbert-base-uncased-finetuned-sst-2-english 68.21 48.00
Pooling ERM (linear) bert-base-multilingual-uncased-sentiment 72.14 53.33
@@
gray!80blackPooling ERM (MLP) bert-base-multilingual-uncased-sentiment 73.01 53.33
@@
gray!80blackDI-ERM (Domain2Vec) bert-base-multilingual-uncased-sentiment 73.19 54.67
@@ D.3 Image classification across styles
For all the experiment we use the text prompt in Figure˜5 as input to DistillBERT.
@@
We use different models to extract the visual features. Then train a logistic regression on top of these features using Scikit-learn with lbfgs solver, the results are deterministic. Section˜D.3 summarizes the results.
Across most domain shifts, our proposed DI-ERM method consistently outperforms standard pooling ERM, highlighting the advantage of incorporating domain-specific information into the representation.
Notably, we observe that the performance gains from DI-ERM tend to diminish as model capacity increases. For the largest models (e.g., CLIP ViT-L/14 and DINOv2 ViT-L/14), the improvement is marginal or saturates. This trend is also observed by various empirical works, e.g. Cho et al. (2023).
Table 9: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling ERM in most configurations, particularly for mid-sized models.
@@
CLIP: vitb32 Pooling ERM (linear) 86.97 99.58 95.90 97.48 94.98
@@
gray!80blackCLIP: vitb16 Pooling ERM (linear) 90.89 99.70 97.51 98.76 96.70
@@
gray!80blackCLIP: vitl14 Pooling ERM (linear) 95.42 99.94 99.22 99.79 98.59
@@
DINOv2: vits14 Pooling ERM (linear) 79.82 85.81 93.55 91.34 87.63
@@
gray!80blackDINOv2: vitb14 Pooling ERM (linear) 87.27 95.45 97.66 94.67 93.76
@@
gray!80blackDINOv2: vitl14 Pooling ERM (linear) 92.29 96.41 98.14 97.48 96.08
@@
DINOv3: vits16 Pooling ERM (linear) 85.14 94.25 94.34 92.92 91.67
@@
gray!80blackDINOv3: vitb16 Pooling ERM (linear) 93.61 79.58 94.58 97.53 91.33
@@
gray!80blackDINOv3: vitl16 Pooling ERM (linear) 96.18 92.28 94.04 99.1 95.40
DI-ERM (linear) 95.80 95.75 94.14 99.1 96.20

### v2

@@ C.4 Proof of Theorem 3
$\mathbb{E}_{X,D,M}\left[\gamma(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]\leq R^{*}_{\mathrm{DI}}-R^{*}_{\mathrm{full}}\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$
@@
$f^{*}_{\mathrm{full}}(X,D)=f^{*}_{\mathrm{DI}}(X,M)\quad\text{ almost surely w.r.t. }P_{XMD}\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DI}}.$
@@
$R^{*}_{\mathrm{DI}}-R^{*}_{\mathrm{full}}$ $=\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{DI}}(X,D)}\right]-\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{full}}(X,M)}\right]$
$=\mathbb{E}_{X,D,M}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(X,M)\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(X,D)\right)|X,D,M\right]$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(x,d)|X=x,D=d,M=m\right)$
$=\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,m)|X=x,D=d\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(x,d)|X=x,D=d\right)$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(x,d)|X=x,D=d,M=m\right)$
@@
$f^{*}_{\mathrm{full}}(x,d)=f^{*}_{\mathrm{DI}}(x,m)\quad\text{ almost surely w.r.t. }P_{XMD}.\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DI}}$
C.5 Proof of Proposition 2
Proposition  For function classes $\{\mathcal{F}_{k}\}$ and $\{\mathcal{G}_{k}\}$ satisfying Assumption 1. Under UBC,
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\geq R^{*}_{\mathrm{DI},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f),\quad\forall k,$ (6)
@@ C.5 Proof of Proposition 2
$\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R^{*}_{\mathrm{DI},\mathcal{F}_{k}}$ (7)
Proof.
The first line follows from the first assumption of Assumption 1, where every function $g\in\mathcal{G}_{k}$ is also realizable in $\mathcal{F}_{k}$.
Therefore, we have
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq R^{*}_{\mathrm{DI},\mathcal{F}_{k}}\geq R^{*}_{\mathrm{DI}}\quad\forall k$
Take the limit w.r.t. $k$
$\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq R^{*}_{\mathrm{DI}}$
Recall from the universal approximation property of function class $\mathcal{G}_{k}$, we have
$\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}=R^{*}_{\mathrm{pool}}.$
What’s more, under UBC assumption, we have
$R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}.$
Then it must holds that
$\lim_{k\to\infty}R^{*}_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R^{*}_{\mathrm{DI},\mathcal{F}_{k}}$
@@ Appendix D Experimental details
The following subsections follows the same structure as Section 6, while providing additional details and full tables.
@@ D.1 Annotator disagreement
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate the sentence $x$ with the annotator profile $m$ using the text prompt shown in Figure 3.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 10 epochs with learning rate $5\cdot 10^{-5}$. Table 5 reports the results over three trials. Our models consistently outperform prior work, with the best configuration achieving over $90\%$ test accuracy—substantially higher than the previous state-of-the-art reported by Deng et al. [2023].
@@
Table 6 presents the results. DI-ERM consistently outperforms pooling-ERM across different feature extractors.
Table 6: Test accuracy on the sentiment disagreement dataset (frozen feature extractor). DI-ERM consistently outperforms pooling-ERM, and in some settings surpasses the prior state-of-the-art of Deng et al. [2023]. We highlight the best performance reported by Deng et al. [2023] (69.77) and our highest score (83.41). $\dagger$: Checkpoints used in Deng et al. [2023] were not publicly specified.
@@
AnnEmb [Deng et al., 2023] BERT 64.61
@@
pooling-ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 45.85
@@
pooling-ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 55.07
@@
pooling-ERM (linear) bert-base-multilingual-uncased-sentiment 43.06
@@
pooling-ERM (MLP) bert-base-multilingual-uncased-sentiment 53.90
@@ D.2 Reviewer-specific analysis
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate each review $x$ with reviewer context $m$, represented by 20 randomly selected reviews from the same reviewer, using the text prompt in Figure 4.
We choose nomic-embed-text-v1.5 [Nussbaum et al., 2025], which supports a context window of up to 8012 tokens (we choose 2048 in our experiments), in order to handle the long reviwer context $m$.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 2 epochs with learning rate $10^{-5}$. Table 7 reports the results over three trials. DI-ERM achieves the best performance, outperforming previously reported methods on the WILDS leaderboard (https://wilds.stanford.edu/).
@@
Table 7: Reviewer-specific sentiment analysis. DI-ERM (ours) achieves the highest accuracy, outperforming prior state-of-the-art by Yao et al. [2022].
@@
Pooling-ERM nomic-embed-text 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
D3G nomic-embed 71.8 $\pm$ 0.1 54.7 $\pm$ 0.0
@@
Inspired by Zaheer et al. [2017], Deshmukh et al. [2018], we implement a Domain2Vec-style module to encode reviewer-specific domain information. Given a set of reviews $\{x_{1},x_{2},\dots,x_{n}\}\sim P_{X|D=d}$ written by reviewer $d$, we learn a mapping
@@
Table 8 shows the result.
Table 8: Sentiment classification on Amazon-WILDS with reviewer-specific signals. “Domain2Vec” denotes reviewer encoding based on a learned mean embedding. DI-ERM variants consistently outperform pooling-ERM baselines.
@@
Pooling-ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 67.42 48.00
@@
Pooling-ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 67.59 48.00
@@
DI-ERM (Domain2Vec) distilbert-base-uncased-finetuned-sst-2-english 68.21 48.00
Pooling-ERM (linear) bert-base-multilingual-uncased-sentiment 72.14 53.33
@@
Pooling-ERM (MLP) bert-base-multilingual-uncased-sentiment 73.01 53.33
@@
DI-ERM (Domain2Vec) bert-base-multilingual-uncased-sentiment 73.19 54.67
@@ D.3 Image classification across styles
For all the experiment we use the text prompt in Figure 5 as input to DistillBERT.
@@
We use different models to extract the visual features. Then train a logistic regression on top of these features using Scikit-learn with lbfgs solver, the results are deterministic. Table 9 summarizes the results.
Across most domain shifts, our proposed DI-ERM method consistently outperforms standard pooling-ERM, highlighting the advantage of incorporating domain-specific information into the representation.
Notably, we observe that the performance gains from DI-ERM tend to diminish as model capacity increases. For the largest models (e.g., CLIP ViT-L/14 and DINOv2 ViT-L/14), the improvement is marginal or saturates. This trend is also observed by various empirical works, e.g. Cho et al. [2023].
Table 9: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling-ERM in most configurations, particularly for mid-sized models.
@@
CLIP: vitb32 Pooling-ERM (linear) 86.97 99.58 95.90 97.48 94.98
@@
CLIP: vitb16 Pooling-ERM (linear) 90.89 99.70 97.51 98.76 96.70
@@
CLIP: vitl14 Pooling-ERM (linear) 95.42 99.94 99.22 99.79 98.59
@@
DINOv2: vits14 Pooling-ERM (linear) 79.82 85.81 93.55 91.34 87.63
@@
DINOv2: vitb14 Pooling-ERM (linear) 87.27 95.45 97.66 94.67 93.76
@@
DINOv2: vitl14 Pooling-ERM (linear) 92.29 96.41 98.14 97.48 96.08
@@
DINOv3: vits16 Pooling-ERM (linear) 85.14 94.25 94.34 92.92 91.67
@@
DINOv3: vitb16 Pooling-ERM (linear) 93.61 79.58 94.58 97.53 91.33
@@
DINOv3: vitl16 Pooling-ERM (linear) 96.18 92.28 94.04 99.10 95.40
DI-ERM (linear) 95.80 95.75 94.14 99.10 96.20